# Physics-informed graph neural networks for flow field estimation in carotid arteries

### v1

@@ 3.2.4 Physics-based loss regularisation
Hemodynamics in the carotid arteries can be modelled by the Navier-Stokes equations. In particular, the divergence $\nabla\cdot u$ of the velocity field $u$ describes conservation of mass and is a recommended quality control in 4D flow MRI [7]. Both conservation of mass and momentum can additionally function as training loss regularisation and accuracy metrics, promoting and measuring the conformity with the underlying physics.
In the context of numerical methods for partial differential equations (PDE) as well as machine learning, spatial discretisation of such differential operators is predominantly done via the finite element method [32] or graph exterior calculus [33]. Both require a mesh of the volumetric, spatial domain consisting of simplices (e.g. tetrahedra) which poses two problems for our application. Firstly, velocity field measurements from 4D flow MRI, which we use as training data, do not come with a volumetric mesh and algorithmic mesh creation is a non-trivial task. Secondly, tetrahedral meshes typically feature a lot more tetrahedra than vertices. Finite-element-discretised operators are sparse matrices containing basis function coefficients defined on the mesh elements, which means their number of non-zero entries scales with the number of tetrahedra. This scaling is reflected in high memory footprint and compute overhead. To circumvent these challenges, we derive a novel, compute-efficient discretisation scheme, which uses neighbourhood queries.
@@
$\nabla\cdot u=\frac{\partial u_{1}}{\partial x_{1}}+\frac{\partial u_{2}}{\partial x_{2}}+\frac{\partial u_{3}}{\partial x_{3}}=0,$
@@
We combine these with $\mathrm{L}^{1}$ loss for the data term in the training loss. We balance the loss terms by multiplying a constant, such that loss term values roughly coincide at model initialisation.
@@ 3.3 Evaluation metrics
[…]
which measures the similarity between two vector fields. Furthermore, we use mean cosine similarity of two vector field
[…]
@@ 3.3.1 Regularised optimal transport distance
Both approximation disparity and cosine similarity are useful metrics if point-wise comparison between velocity fields is possible. Beyond this case, we can only rely on comparison between local neighbourhoods of velocity vectors. A metric measuring the difference between local sets of vectors may be more robust to extreme outliers (as in noisy measurements) and can be used to compare discrete flow fields where the spatial positions do not coincide. To this end, we propose to use a regularised optimal transport distance, measuring the divergence between distributions of vectors. We compute the metric using an efficient implementation of the Sinkhorn divergence [34] for approximating optimal transport cost, which is effectively a relaxation of the Wasserstein distance. Our metric takes into account the spatial position as well as direction and magnitude of two velocity vector fields.
@@
Figure 4: Results of neural network predictions (“ML”) by PointNet++ compared with 4D flow MRI in the left and right carotid artery of subjects in the test split. We visualise the velocity field via 3D streamlines and render a projection image. Shown are examples of relatively a) good, b) average and c) poor performance. Additionally, we d) show an example where the ground truth is noisy, yet the model predicts a visually sound velocity field.
4 Experiments and results
All neural networks were implemented in Python using PyTorch [35] and PyTorch Geometric [36], had approximately 1 million trainable parameters and were trained on an NVIDIA A100 (40 GB) GPU for 1000 epochs using Adam optimiser (learning rate $8\cdot 10^{-4}$ with exponential decay of 0.9955) with gradient clipping. Training PointNet++ took 11:13 s per epochs compared to 19:20 s for the equivariant SE-PointNet++. Runtime overhead of the physics-informed loss regularisation was negligible. The regularised Wasserstein distance computation (see Section 3.3.1) was implemented via GeomLoss [34]. We evaluated our neural networks via cross-validation. Our implementation is publicly available.111github.com/sukjulian (coming soon)
We partitioned the 282 subject-specific 4D flow MRI measurements ten-fold into 254 training and 28 evaluation samples, so that all subjects appeared in exactly one evaluation split (excluding two remainders). Left and right carotid artery of the same subject were considered separate samples, but we made sure that both were contained in the same fold.
4.1 Velocity field estimation in 4D flow MRI
4.1.1 PointNet++
In Table 1 we compare test split predictions $u^{p}$ of PointNet++ to the corresponding 4D flow MRI measurements $\bar{u}^{p}$ via Wasserstein distance (see Section 3.3.1), approximation disparity and cosine similarity across all test-split subjects simultaneously. Since some of the 4D flow MRI measurements were noisy (compare Fig. 4 d)) we assess correspondence to the Navier-Stokes equations via continuity (3) and momentum residual (4). PointNet++ achieves good correspondence to 4D flow MRI, indicated by cosine similarity close to one. Fig. 4 visualises the reference velocity field computed via 4D flow MRI in the left and right carotid artery of four subjects from the combined cross-validation test splits. Comparison with the estimated velocity fields by PointNet++ shows good overall qualitative agreement to the reference. Even for the case where the 4D flow MRI velocity field is noisy (Fig. 4 d)) PointNet++ produces a plausible estimate.
4.1.2 SE-PointNet++
SE-PointNet++ obtained slightly higher approximation disparity than PointNet++, but we did not find the difference to be statistically significant ($p<0.05$) in a one-way ANOVA test. SE-PointNet++ estimates had better conservation of mass and momentum than PointNet++ which is also supported by visual inspection. Fig. 5 shows a comparison between velocity field estimates by PointNet++ and by SE-PointNet++ and thus the influence of group equivariance. We found that SE-PointNet++ produced slightly smoother estimates while skipping over noise present in the PointNet++ estimate, e.g. in the ICA of the right carotid artery in Fig. 5 a).
4.1.3 Physics-informed extensions
Using physics-informed loss regularisation (Section 3.2.4), we found that continuity and momentum residuals decreased even further. Additionally, we observed a slight, albeit statistically insignificant, decrease in approximation disparity for both neural networks. We conclude that physics-based loss regularisation is independently beneficial for accuracy, i.e., does not compete with the training objective of fitting the data. SE-PointNet++${}^{\text{PIGN}}$ performs best w.r.t. correspondence to the continuity and momentum residual. In Fig. 6 we compare velocity field estimates of SE-PointNet++ trained with and without physics-informed loss regularisation (SE-PointNet++${}^{\text{PIGN}}$) based on the Navier-Stokes equations, as described in Section 3.2.4. We found that this further smoothed the flow field, e.g. in the ECA of the right carotid artery in Fig. 6 a).
@@ 4.1.3 Physics-informed extensions
Table 1: Quantitative evaluation with respect to the 4D flow MRI measurements and flow physics (Navier-Stokes), evaluated via 10-fold cross-validation. We show mean $\pm$ standard deviation (across subjects) of the statistics. Since there is no point-to-point correspondence between black blood MRI and 4D flow MRI, Approx. disp. and Cos. similarity are omitted. We denote in bold superior performance where it significantly ($p<0.05$) differs according to a one-way ANOVA test.
[…]
@@
Figure 6: Effect of physics-informed training (“PIGN”) for the example of SE-PointNet++. We show the left and right carotid artery of subjects in the test split and visualise the velocity field via streamlines.
@@ 4.1.4 Ablation studies
We created three ablated models to investigate the task-specific drivers of performance in PointNet++. Firstly, we trained a simple MLP that maps the input features of each point to a velocity vector, without local or global interaction between points. Secondly, we trained an “empty” PointNet++ without geometric input features, i.e. point-wise description of relative geometry (see Section 3.2.1), which learned only from message passing between points (and boundary conditions). Thirdly, we trained a PointNet++ with geometric input features but without inflow conditioning (see Section 3.2.1). Performance metrics are given in Table 1. We found that all three variant achieves significantly worse approximation disparity than the baseline PointNet++. MLP achieved lower approximation disparity than the “empty” PointNet++ pointing to the importance of the geometric input features. Training time per epoch was 11:02 s for the MLP and thus runtime was comparable to PointNet++. Furthermore, we found that without inflow conditioning, PointNet++ significantly dropped in accuracy. This is expressed in higher approximation disparity but not lower cosine similarity, pointing towards magnitude rather than direction of predicted estimates.
@@ 4.2 Generalisation to black-blood MRI
We investigate our neural networks’ capabilities to generalise between vascular models obtained using different imaging protocols, by training the neural networks on the 4D flow MRI dataset and letting them estimate velocity fields in 3D models of the same artery but obtained via black-blood MRI. We pass the inflow boundary conditions extracted from the inlet region of the 4D flow MRI measurements (see Section 3.2.1) to the machine learning model, simulating in-vivo measurements via e.g. ultrasound. Note that black blood MRI in itself does not allow for velocity measurements, since it is not a temporal imaging method and does not measure blood flow or contains flow-encoding sequence parts. Since black-blood MRI by itself does not allow for velocity measurements, we quantify our neural networks’ accuracy by Wasserstein distance (see Section 3.3.1) to the patient-specific 4D flow measurements as well as continuity and momentum residual (see Section 3.2.4). Table 1 lists the results. PointNet++ achieves the lowest mean Wasserstein distance, but the differences are not significant. As above, SE-PointNet++${}^{\text{PIGN}}$ achieves the lowest continuity and momentum residuals. These are slightly higher than for the 4D flow MRI data. Nevertheless, $L_{\text{continuity}}$ is substantially lower in flow field estimates produced by SE-PointNet++${}^{\text{PIGN}}$ than in those of PointNet++, SE-PointNet++ and PointNet++${}^{\text{PIGN}}$. Fig. 7 shows a comparison between ground truth 4D flow MRI and machine learning estimates (denoted “black blood & ML”) in 3D carotid artery models of subjects from the test split obtained via black blood MRI. We use SE-PointNet++ for this, since it, unlike PointNet++, does not require alignment of the black blood MRI geometry with the 4D flow MRI due to its roto-translation equivariance. This is to simulate a realistic scenario where 4D flow MRI would not be available. We find good agreement of the machine learning model with the ground truth, e.g. in the left carotid artery in Fig. 7 a) and the right carotid artery in Fig. 7 b). For several cases the machine learning model underestimates the magnitude of the flow (Fig. 7 d)). Nevertheless, we find overall good qualitative agreement in the estimated and measured flow fields.
Figure 7: Generalisation to black-blood MRI in the left and right carotid bifurcation of subjects from the test split. We compare 4D flow MRI with application of a pre-trained machine learning model on geometries obtained via black blood MRI (denoted “black blood & ML”). Note that black blood MRI itself does not provide measurements of the flow field.
@@ 5 Discussion and conclusion
In this work, we present a novel deep learning approach for modelling hemodynamics by generating velocity fields for unseen subject-specific carotid artery geometries. We build upon the PointNet++ [24] architecture and infuse it with steerable equivariant [1] MLPs. We show how this symmetry consideration implicitly leads to stronger adherence to the underlying flow physics. Our neural network architecture allows for global context aggregation via cascading of pooling and interpolation layers and bypasses the need for message passing for each point in the original resolution, which would be computationally prohibitive. Thus, we are able to effectively learn features mapped to the original point cloud at low computational cost. We complement our deep learning approach by including domain knowledge, specifically about the differential structure of the velocity field, by regularising the training loss with discretised residuals based on the Navier-Stokes equations. To this end, we derive a computationally efficient, mesh-free discretisation scheme whose complexity is linear in the number of points. Ultimately, we demonstrate how our method can learn to estimate velocity fields in unseen subjects after being trained exclusively on 4D flow MRI data.
@@
The approach we propose addresses some important limitations of the state-of-the-art methods in machine-learning-based flow field estimation (Section 2). Compared to single-case instance optimisation methods for fluid flow estimation [15, 16, 17], our approach is able to generalise beyond the training data while still respecting the governing PDEs. Compared to previous works on multiple-case generalising feed-forward methods [19, 20, 21, 22, 11, 14], our model gains independence of the variability of the CFD-based training data, because we train it on in-vivo flow field measurements from 4D flow MRI.
@@
Although we used 4D flow MRI training data to train and validate our method, we also demonstrate that we can train machine learning models on 4D flow MRI velocity fields and transfer the learned relations to vascular models obtained from anatomical black blood MRI images. This could be valuable in practice, as anatomical imaging modalities might be more widely available than 4D flow MRI and 3D vascular models can be obtained from these modalities automatically, e.g., in black blood MRI [30], CT [8], and 3D ultrasound. We envision this approach to be used in clinical practice for fast estimation of patient-specific blood flow, but backed up by 4D flow MRI for cases in which complications are suspected. This would still require personalised boundary conditions, which could be non-invasively obtained using, e.g., Doppler ultrasound. Note however, that even though these geometries are of the same patient, different imaging modalities produce slightly different geometries which is a source of error in our method.
@@
In 4D flow MRI, velocity cannot be accurately computed very close to the artery wall which results in high levels of noise. In our experiments, even though the training data was noisy in some cases, our neural networks did not reproduce the noise but implicitly smoothed over it in their estimations. We attribute this to the fact that the noise cannot be explained by the input anatomy in a meaningful way and our neural networks treat it as constant error term. This has potential implications for the estimation of wall shear stress which requires noise-free velocity estimates close to the artery wall.
@@
Since our method is a data-driven approach, the quality of the results largely depends on the extent to which the training data represents the data distribution during inference. In contrast to comparable studies using synthetic flow data generated by CFD [19, 20, 21, 22, 11, 14], we have access to a small dataset in this work. This poses no major challenge because the carotid bifurcation across different subjects in our data were of a relatively similar shape. However, it would be interesting to see what happens for larger variations in the data. Besides stenotic cardiovascular disease (CVD), aneurysmatic CVD, in the form of intracranial and abdominal aneurysms, is also a dangerous disease which depends on blood flow. Thus, aneurysmatic CVD also provides potential applications of our methods in determination of hemodynamic quantities in these kinds of diseases. Even though we focussed on the carotid arteries in this work, we expect our approach to apply to other arteries as well, as long as there is a sufficiently representative dataset relative to the artery complexity (e.g. stenoses, aneurysms or increased tortuosity). What is more, the quantitative evaluation in our experiments is held back by the measurement noise present in some of the test cases since it obfuscates the expressiveness of the evaluation metrics. We address this by performing statistical significance tests to discern differences in the numerical results. Lastly, our analysis in this work is limited to time-averaged instead of pulsatile hemodynamics estimation, even though 4D flow MRI does enable transient quantification of blood flow.

### v2

@@ 3.2.4 Physics-based loss regularisation
Hemodynamics in the carotid arteries can be modelled by the Navier-Stokes equations. In particular, the divergence $\nabla\cdot u$ of the velocity field $u$ describes conservation of mass and is a recommended quality control in 4D flow MRI Bissell et al. (2023). Both conservation of mass and momentum can additionally function as training loss regularisation and accuracy metrics, promoting and measuring the conformity with the underlying physics.
In the context of numerical methods for partial differential equations (PDE) as well as machine learning, spatial discretisation of such differential operators is predominantly done via the finite element method Gao et al. (2022) or graph exterior calculus Shukla et al. (2022). Both require a mesh of the volumetric, spatial domain consisting of simplices (e.g. tetrahedra) which poses two problems for our application. Firstly, velocity field measurements from 4D flow MRI, which we use as training data, do not come with a volumetric mesh and algorithmic mesh creation is a non-trivial task. Secondly, tetrahedral meshes typically feature a lot more tetrahedra than vertices. Finite-element-discretised operators are sparse matrices containing basis function coefficients defined on the mesh elements, which means their number of non-zero entries scales with the number of tetrahedra. This scaling is reflected in high memory footprint and compute overhead. To circumvent these challenges, we derive a novel, compute-efficient discretisation scheme, which uses neighbourhood queries.
@@
$\nabla\cdot u=\frac{\partial u_{1}}{\partial x_{1}}+\frac{\partial u_{2}}{\partial x_{2}}+\frac{\partial u_{3}}{\partial x_{3}}=0.$
@@
We combine these with $\mathrm{L}^{1}$ loss for the data term in the training loss. We balance the loss terms by multiplying a constantbased on observed training convergence.
@@ 3.3 Evaluation metrics
$\text{Approx. disp.}\colon\sqrt{\sum_{p\in\mathcal{P}_{0}}\|\bar{u}^{p}-u^{p}\|_{2}^{2}/\sum_{p\in\mathcal{P}_{0}}\|\bar{u}^{p}\|_{2}^{2}}$
which measures the similarity between two vector fields. Furthermore, we use mean cosine similarity of two vector fields
$\text{Cos. similarity}\colon\operatorname*{mean}_{p\in\mathcal{P}_{0}}\cos(\angle\bar{u}^{p},u^{p})$
@@ 3.3.1 Regularised optimal transport distance
Both approximation disparity and cosine similarity are useful metrics if point-wise comparison between velocity fields is possible. Beyond this case, we can only rely on comparison between local neighbourhoods of velocity vectors after spatial alignment (registration) of the point clouds. A metric measuring the difference between local sets of vectors may be more robust to extreme outliers (as in noisy measurements) and can be used to compare discrete flow fields where the spatial positions do not coincide. To this end, we propose to use a regularised optimal transport distance, measuring the divergence between distributions of vectors. We compute the metric using an efficient implementation of the Sinkhorn divergence Feydy et al. (2019) for approximating optimal transport cost, which is effectively a relaxation of the Wasserstein distance. Local neighbourhoods of points are constructed by first aligning both point clouds via Procrustes analysis and then including ground truth positions within a certain radius around the positions of model prediction. Our metric takes into account the spatial position as well as direction and magnitude of two velocity vector fields.
@@
Figure 4: Results of neural network predictions (“ML") by PointNet++ compared with 4D flow MRI in the left and right carotid artery of subjects in the test split. We visualise the velocity field via 3D streamlines and render a projection image. Shown are examples of relatively a) good, b) average and c) poor performance. Additionally, we d) show an example where the ground truth is noisy, yet the model predicts a visually sound velocity field.
@@
Table 1: Quantitative evaluation with respect to the 4D flow MRI measurements and flow physics (Navier-Stokes), evaluated via 10-fold cross-validation. We show mean $\pm$ standard deviation (across subjects) of the statistics. Since there is no point-to-point correspondence between black blood MRI and 4D flow MRI, Approx. disp. and Cos. similarity are omitted. We denote in bold superior performance where it significantly ($p<0.05$) differs according to a one-way ANOVA test: per metric, we picked the model with the lowest mean and tested significance compared to all other models. We consider all models whose metric does not differ significantly as equal.
Imaging modality Neural network Wasserst. dist. $\cdot\texttt{1e3}$ $\downarrow$ Approx. disp. $\downarrow$ Cos. similarity $\uparrow$ $L_{\mathrm{continuity}}$ $\downarrow$ $L_{\mathrm{momentum}}\cdot\texttt{1e-3}$ $\downarrow$
@@
4 Experiments and results
All neural networks were implemented in Python using PyTorch Paszke et al. (2019) and PyTorch Geometric Fey and Lenssen (2019), had approximately 1 million trainable parameters and were trained on an NVIDIA A100 (40 GB) GPU for 1000 epochs using Adam optimiser (learning rate $8\cdot 10^{-4}$ with exponential decay of 0.9955) with gradient clipping. Training PointNet++ took 11:13 s per epochs compared to 19:20 s for the equivariant SE-PointNet++. Runtime overhead of the physics-informed loss regularisation was negligible. The regularised Wasserstein distance computation (see Section 3.3.1) was implemented via GeomLoss Feydy et al. (2019). We evaluated our neural networks via cross-validation. Our implementation is publicly available.111github.com/sukjulian/physics-informed-gnn
We partitioned the 282 subject-specific 4D flow MRI measurements 10-fold into 254 training and 28 evaluation samples, so that all subjects appeared in exactly one evaluation split (excluding two remainders). Left and right carotid artery of the same subject were considered separate samples, but we made sure that both were contained in the same fold. For training curves, see Appendix B.
4.1 Velocity field estimation in 4D flow MRI
4.1.1 PointNet++
In Table 1 we compare test split predictions $u^{p}$ of PointNet++ to the corresponding 4D flow MRI measurements $\bar{u}^{p}$ via Wasserstein distance (see Section 3.3.1), approximation disparity and cosine similarity across all test-split subjects simultaneously. [^2] For explanation of the high standard deviation in Wasserstein distance see Appendix A. Since some of the 4D flow MRI measurements were noisy (compare Fig. 4 d)) we assess correspondence to the Navier-Stokes equations via continuity (3) and momentum residual (4). PointNet++ achieves good correspondence to 4D flow MRI, indicated by cosine similarity close to one. Fig. 4 visualises the reference velocity field computed via 4D flow MRI in the left and right carotid artery of four subjects from the combined cross-validation test splits. Comparison with the estimated velocity fields by PointNet++ shows good overall qualitative agreement to the reference. Even for the case where the 4D flow MRI velocity field is noisy (Fig. 4 d)) PointNet++ produces a plausible estimate.
4.1.2 SE-PointNet++
SE-PointNet++ obtained the same approximation disparity as PointNet with no statistically significant ($p<0.05$) difference in a one-way ANOVA test. SE-PointNet++ estimates had better conservation of mass and momentum than PointNet++ which is also supported by visual inspection. Fig. 5 shows a comparison between velocity field estimates by PointNet++ and by SE-PointNet++ and thus the influence of group equivariance. We found that SE-PointNet++ produced slightly smoother estimates while skipping over noise present in the PointNet++ estimate, e.g. in the ICA of the right carotid artery in Fig. 5 a).
4.1.3 Physics-informed extensions
[…]
@@ 4.1.3 Physics-informed extensions
Figure 6: Influence of supervised loss in the presence of physics-informed loss regularisation. We trained PointNet++ for different supervised loss term weighting factors and computed Approx. disp. w.r.t. the 4D flow MRI data as well as $L_{\mathrm{continuity}}$ and $L_{\mathrm{momentum}}$ (corresponding to the physics-informed loss regularisation). In each case, we performed 10-fold cross-validation resulting in distributions which we visualise via boxplot.
@@ 4.1.4 Ablation studies
We created three ablated models to investigate the task-specific drivers of performance in PointNet++. Firstly, we trained a simple MLP that maps the input features of each point to a velocity vector, without local or global interaction between points. In this ablation, all points are treated independently only based on their (input) features and the model does not have spatial context. Secondly, we trained an “empty" PointNet++ without geometric input features, i.e. point-wise description of relative geometry (see Section 3.2.1), which learned only from message passing between points (and boundary conditions). Thirdly, we trained a PointNet++ with geometric input features but without inflow conditioning (see Section 3.2.1). Performance metrics are given in Table 1. We found that all three variant achieved significantly worse approximation disparity than the baseline PointNet++. MLP and the “empty" PointNet++ did not have statistically significant difference in approximation disparity, pointing to equal importance of geometric input features and spatial context. Training time per epoch was 11:02 s for the MLP and thus runtime was comparable to PointNet++. Furthermore, we found that without inflow conditioning, PointNet++ significantly dropped in accuracy. This is expressed in higher approximation disparity but not lower cosine similarity, pointing towards magnitude rather than direction of predicted estimates.
Figure 7: Effect of physics-informed training (“PIGN") for the example of SE-PointNet++. We show the left and right carotid artery of subjects in the test split and visualise the velocity field via streamlines.
4.1.5 Subject-specific physics-informed graph neural network(s)
As additional reference for comparison besides 4D flow MRI, we trained – for each subject – a physics-informed graph neural network leveraging our physics-informed loss regularisation. We used PointNet++ for fast training and weighted the data term by 4, the continuity term by 0.01 and the momentum term by 1e-5 based on observed training convergence. Compared to the 4D flow MRI, the resulting model had mean ($\pm$ standard deviation) Approx. disp. of 0.29 ($\pm$ 0.14) across subjects. Physics-informed residuals $L_{\mathrm{continuity}}$ and $L_{\mathrm{momentum}}\cdot\texttt{1e-3}$ were 1.7 ($\pm$ 0.6) and 2.8 ($\pm$ 1.4), respectively. In Table 2 we present approximation disparity of all models (trained on 4D flow MRI) compared to the physics-informed graph neural network. We found the same relative accuracy as in Table 1 with the only difference that PointNet++ was slightly less accurate than its physics-informed counterpart here.
@@ 4.2 Generalisation to black-blood MRI
[…]
Table 2: Comparison to subject-specific physics-informed graph neural network via 10-fold cross-validation. We show mean $\pm$ standard deviation (across subjects). We denote in bold superior performance where it significantly ($p<0.05$) differs according to a one-way ANOVA test: per metric, we picked the model with the lowest mean and tested significance compared to all other models. We consider all models whose metric does not differ significantly as equal.
Neural network Approx. disp. $\downarrow$
MLP 0.54 $\pm$ 0.19
PointNet++ 0.50 $\pm$ 0.18
SE-PointNet++ 0.49 $\pm$ 0.18
[…]
@@ 5 Discussion and conclusion
In this work, we present a novel deep learning approach for modelling hemodynamics by generating velocity fields for unseen subject-specific carotid artery geometries. We build upon the PointNet++ Qi et al. (2017) architecture and infuse it with steerable equivariant Brandstetter et al. (2022) MLPs. We show how this symmetry consideration implicitly leads to stronger adherence to the underlying flow physics. Our neural network architecture allows for global context aggregation via cascading of pooling and interpolation layers and bypasses the need for message passing for each point in the original resolution, which would be computationally prohibitive. Thus, we are able to effectively learn features mapped to the original point cloud at low computational cost. We complement our deep learning approach by including domain knowledge, specifically about the differential structure of the velocity field, by regularising the training loss with discretised residuals based on the Navier-Stokes equations. To this end, we derive a computationally efficient, mesh-free discretisation scheme whose complexity is linear in the number of points. Ultimately, we demonstrate how our method can learn to estimate velocity fields in unseen subjects after being trained exclusively on 4D flow MRI data.
@@
The approach we propose addresses some important limitations of the state-of-the-art methods in machine-learning-based flow field estimation (Section 2). Compared to single-case instance optimisation methods for fluid flow estimation Raissi et al. (2020); Fathi et al. (2020); Kontogiannis and Juniper (2022), our approach is able to generalise beyond the training data while still respecting the governing PDEs. Compared to previous works on multiple-case generalising feed-forward methods Liang et al. (2020); Wang et al. (2023); Maul et al. (2023); Li et al. (2021); Zhang et al. (2023); Suk et al. (2023), our model gains independence of the variability of the CFD-based training data, because we train it on in-vivo flow field measurements from 4D flow MRI.
Figure 8: Generalisation to black-blood MRI in the left and right carotid bifurcation of subjects from the test split. We compare 4D flow MRI with application of a pre-trained machine learning model on geometries obtained via black blood MRI (denoted “black blood & ML"). Note that black blood MRI itself does not provide measurements of the flow field.
@@
Although we used 4D flow MRI training data to train and validate our method, we also demonstrate that we can train machine learning models on 4D flow MRI velocity fields and transfer the learned relations to vascular models obtained from anatomical black blood MRI images. This could be valuable in practice, as anatomical imaging modalities might be more widely available than 4D flow MRI and 3D vascular models can be obtained from these modalities automatically, e.g., in black blood MRI Alblas et al. (2021), CT Li et al. (2024) and 3D ultrasound. We envision this approach to be used in clinical practice for fast estimation of patient-specific blood flow, but backed up by 4D flow MRI for cases in which complications are suspected. This would still require personalised boundary conditions, which could be non-invasively obtained using, e.g., Doppler ultrasound. [^4] Doppler ultrasound comes with its own set of drawbacks such as its restriction to 1D and its dependency on the measurement angle and user. Note however, that even though these geometries are of the same patient, different imaging modalities produce slightly different geometries which is a source of error in our method.
@@
In 4D flow MRI, velocity cannot be accurately computed very close to the artery wall which results in high levels of noise. In our experiments, even though the training data was noisy in some cases, our neural networks did not reproduce the noise but implicitly smoothed over it in their estimations. We attribute this to the fact that the noise cannot be explained by the input anatomy in a meaningful way and our neural networks treat it as constant error term. It would be interesting to investigate to what extent the estimated velocity field could be smoothed using a similar strategy as our physics-informed loss regularisation (Section 3.2.4) and if this could enable extraction of wall shear stress which requires noise-free velocity estimates close to the artery wall.
We found that our $\mathrm{E(3)}$-equivariant neural network a priori produced estimates with significantly lower Navier-Stokes residuals compared to the vanilla baseline. This can be partly explained by the inherent symmetries present in fluid flow. However, this phenomenon should be systematically studied and is an interesting avenue for future work.
@@
Since our method is a data-driven approach, the quality of the results largely depends on the extent to which the training data represents the data distribution during inference. For instance, a model trained on arteries of a certain topology (number of outles, bifurcations, etc.) is unlikely to generalise to substantially different topologies. This might also apply to scans with substantially different field of view (up- or downstream) dependent on the specific representation the model has learned (e.g. dependence on absolute coordinates). In contrast to comparable studies using synthetic flow data generated by CFD Liang et al. (2020); Wang et al. (2023); Maul et al. (2023); Li et al. (2021); Zhang et al. (2023); Suk et al. (2023), we have access to a small dataset in this work. This poses no major challenge because the carotid bifurcation across different subjects in our data are of a relatively similar shape. However, it would be interesting to see what happens for larger variations in the data, both geometric and regarding flow data. This would especially be relevant for large datasets that are collaboratively sourced from multiple institutes: 4D flow MRI can vary significantly across methodologies and hardware. In such a scenario, velocity fields should be canonicalised for optimal model performance. In principle, we could also including well-tuned CFD simulations in our training data to complement the 4D flow MRI, as the former can also be represented as point clouds of velocity vectors. Such a hybrid-data approach is an interesting avenue for future research. As data-drive approach, our method inherits the dependency on acquisition settings and image quality from the 4D flow MRI on which it is trained. Just like 4D flow MRI, comparability of the generated velocity fields with those acquired under different conditions by different clinicians is critical for potential clinical translation. Testing this mechanism with fixed patients would be crucial insight into the robustness of our method and is thus important future research. Besides stenotic cardiovascular disease (CVD), aneurysmatic CVD, in the form of intracranial and abdominal aneurysms, is also a dangerous disease which depends on blood flow. Thus, aneurysmatic CVD also provides potential applications of our methods in determination of hemodynamic quantities in these kinds of diseases. Even though we focussed on the carotid arteries in this work, we expect our approach to apply to other arteries as well, as long as there is a sufficiently representative dataset relative to the artery complexity (e.g. stenoses, aneurysms or increased tortuosity). What is more, the quantitative evaluation in our experiments is held back by the measurement noise present in some of the test cases since it obfuscates the expressiveness of the evaluation metrics. We address this by performing statistical significance tests to discern differences in the numerical results. Lastly, our analysis in this work is limited to single time frame instead of pulsatile hemodynamics estimation, even though 4D flow MRI does enable transient quantification of blood flow. We recognise that single time frame analysis might obfuscate turbulence effects. Especially in the context of stenotic CVD, turbulence plays a prominent role, modelling which is usually performed in transient flow. Incorporating an explicit turbulence model in our physics-informed loss regularisation is an interesting avenue for future research.
@@
Acknowledgement
We would like to thank José Iglesias Martínez for his help in the form of discussions about optimal transport and Sinkhorn divergence. This work is funded in part by the 4TU Precision Medicine programme supported by High Tech for a Sustainable Future, a framework commissioned by the four Universities of Technology of the Netherlands. Jelmer M. Wolterink was supported by the NWO domain Applied and Engineering Sciences Veni grant (18192).
Appendix
Appendix A Standard deviation in Wasserstein distance
See Fig. 9.
Figure 9: Distribution of Wasserstein distance for MLP compared to 4D flow MRI. The orange line denotes the median (Wasserst. dist. $\cdot\texttt{1e3}$ = 7.6). Outliers correspond to 4D flow MRI data with very low signal-to-noise ratio (compare Fig. 4 d)). We did not exclude these outliers to simulate a realistic scenario.
Appendix B Training convergence
See Fig. 10.
Figure 10: Training curves and evaluation metrics progression during training. We show mean Approx. disp., $L_{\mathrm{continuity}}$ and $L_{\mathrm{continuity}}$ over the cross-validation folds. Note that we did not have held-out validation splits for this experiment.